Steam at 100° C is added to ice at 0° C. Find the amount of ice melted and the final temperature when the mass of steam is 7.00 g and the mass of ice is 50.0 g.
Answer #1

DO 202 oc Mass of ice - Given: Temperature of steam = 100°c. Temperature of ice Mass of steam = 7.00g 50.0g Solution: We knowAz = [[7.00)x103)/4186)/100-0) 2930.2 J = 2.930 KI Therefore, Total heat pumided by steam to system - Q₂ + Q = (15.820 + 2.9315.820 - 0.0293 Tft 2.93 Gold = - Quot 16:650 +0.20937, +415.820 +0.0293[T. 0:2386 To = 2:1 => Tp = 8.8°C) (final Temperature